Crane Corp. is a fast-growing company whose management expects it to grow at a rate of 25 percent over the next two years and then to slow to a growth rate of 17 percent for the following three years. If the last dividend paid by the company was $2.15.

What is the dividend for the 1st year? (Round answer to 3 decimal places, e.g. 15.250.)

What is the dividend for the 2nd year? (Round answer to 3 decimal places, e.g. 15.250.)

What is the dividend for the 3rd year? (Round answer to 3 decimal places, e.g. 15.250.)
What is the dividend for the 4th year? (Round answer to 3 decimal places, e.g. 15.250.)
What is the dividend for the 5th year? (Round answer to 3 decimal places, e.g. 15.250.)

Compute the present value of these dividends if the required rate of return is 14 percent. (Round intermediate calculations and final answer to 2 decimal places, e.g. 15.25.)
Answer #1

D0 = 2.15

D1 = 2.15 * 1.25 = 2.688

D2 = 2.15 * 1.25 * 1.25 = 3.359

D3 = 2.15 * 1.25 * 1.25 * 1.17 = 3.930

D4 = 2.15 * 1.25 * 1.25 * 1.17 * 1.17 = 4.599

D5 = 2.15 * 1.25 * 1.25 * 1.17 * 1.17 * 1.17 = 5.380

Present value = 2.688/ (1 + 14%)1 + 3.359/ (1 + 14%)2 + 3.930/ (1 + 14%)3 + 4.599/ (1 + 14%)4 + 5.380/ (1 + 14%)5

Present value = 2.36 + 2.58 + 2.65 + 2.72 + 2.79

Present value = 13.113

Present value = 13.11